OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (anonymous):

all the lines have slope \(-6\) so you have to find the \(y\) intercept to see if it is -15, -20, 15, or 20

OpenStudy (anonymous):

if you have \(y=-6x+b\) and since \((-2,-3)\) is on the graph you know that if \(x=-2\) then \(y=-3\) so you can put \[-3=-6\times (-2)+b\] and solve for \(b\)

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

i got C ?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

yeah i made a typo there, i dont mean to be a pain so lets just solve it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[-3=-6\times (-2)+b\] solve for \(b\) \[-3=12+b\] subtract 12 from both sides and get \[-15=b\]

OpenStudy (anonymous):

so the line is \(y=-6x-15\)

OpenStudy (anonymous):

ohhhhhhhh..... so A

OpenStudy (anonymous):

yah first time i have seen A all week usually C

OpenStudy (anonymous):

i dont quite get these yet.. i have 4 more.. would u mind?

OpenStudy (anonymous):

sure what is the next one?

OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (anonymous):

ok very similar to the last one all your choices have the same slope, namely \(1\) so you only need to find the \(y\) intercept in this case if \(x=6\) you have \(y=-2\) so you can solve \[-2=6+b\] in one easy step subtract \(6\) from both sides

OpenStudy (anonymous):

what do you get this time?

OpenStudy (anonymous):

if you are confused let me know

OpenStudy (anonymous):

so it would be D then?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cool! only 3 more!

OpenStudy (anonymous):

next...

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = 1/2x - 8. y = 1/2x + 7 y = 1/2x + 5 y = 2x + 7 y = 2x + 5

OpenStudy (anonymous):

ok parallel means has the same slope the slope of \(y=\frac{1}{2}x-8\) is evidently \(\frac{1}{2}\) so you can ignore the last two answers because they have slope 2

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

again put \(x=-2,y=6\) and solve \[6=\frac{1}{2}\times (-2)+b\] for \(b\) can you do this or do you need help?

OpenStudy (anonymous):

ill try but what do i do for the first step to get me started?

OpenStudy (anonymous):

be careful with the signs

OpenStudy (anonymous):

first step is to multiply \(\frac{1}{2}\times (-2)\)

OpenStudy (anonymous):

so i got C as my final answer

OpenStudy (anonymous):

we already ruled out C and D right?

OpenStudy (anonymous):

opps meant A

OpenStudy (anonymous):

\[6=\frac{1}{2}\times (-2)+b\] \[6=-1+b\] \[7=b\] right!

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, 3) and perpendicular to y = -1/3x + 7. y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (anonymous):

and cool!

OpenStudy (anonymous):

kk so once again all the slopes are the same

OpenStudy (anonymous):

so guess what? you have to solve \[3=3(-1)+b\] for \(b\) you getting the hang of this?

OpenStudy (anonymous):

let me know what you get for \(b\) again be careful with the signs

OpenStudy (anonymous):

so this one would be B!

OpenStudy (anonymous):

you got it !

OpenStudy (anonymous):

cool only one more!

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, -1) and perpendicular to y = 1/2x + 3. y = -2x - 7 y = -2x +5 y = 2x + 5 y = 2x - 7

OpenStudy (anonymous):

k i will look, this time start and see if you can figure out what you have to solve

OpenStudy (anonymous):

actually before we start, i hope it is clear that the slope of \(y=\frac{1}{2}+3\) is \(\frac{1}{2}\) do you know what the slope of the perpendicular line is?

OpenStudy (anonymous):

oops i meant \(y=\frac{1}{2}x+3\) has slope \(\frac{1}{2}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you know what the slope of the perpendicular line would be? you have two choices in your answers, \(2\) or \(-2\)

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

no actually it is \(-2\) the "negative reciprocal" i .e. flip it and change the sign

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

so that rules out C and D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think that it is B

OpenStudy (anonymous):

lets see point is (-3, -1), slope is \(-2\) so you must solve \[-1=-2\times (-3)+b\] giving \[-1=-6+b\] so \[b=5\] and the equation is \(y=-2x+5\) is that the one you picked?

OpenStudy (anonymous):

lol no... thats ok though! thank you! can i ask you one more? haha

OpenStudy (anonymous):

yea, but you did pick B right?

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-5, 1) with a slope of 3/2 . y - 5 =3/2 (x + 1) y + 1 =3/2 (x - 5) y + 5 = -3/2(x - 1) y - 1 = 3/2(x + 5) and yes i picked B

OpenStudy (anonymous):

y = -2x - 7 y = -2x +5 y = 2x + 5 y = 2x - 7 B is correct

OpenStudy (anonymous):

wait do u mean -2 or 2?

OpenStudy (anonymous):

so is the answer 2x - 7 or -2x-7

OpenStudy (anonymous):

the correct answer was \(y=-2x+5\)

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

oooh nooooo

OpenStudy (anonymous):

let me do it correctly

OpenStudy (anonymous):

i made a mistake, sorry

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

\[-1=-2\times (-3)+b\] \[-1=6+b\] \[-7=b\] \[y=-2x-7\] is the correct answer

OpenStudy (anonymous):

now to your last problem

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

slope is \(m=\frac{3}{2}\) and the point is \((-5,1)\)

OpenStudy (anonymous):

the "point-slope" formula is \[y-y_1=m(x-x_1)\]replace \(m\) by \(\frac{3}{2}\) , replace \(x_1\) by \(-5\) and replace \(y_1\) by \(1\)

OpenStudy (anonymous):

y - 1 = 3/2 ( x + 5) ?

OpenStudy (anonymous):

make a direct substitution with the numbers for the variables and get \[y-1=\frac{3}{2}(x-(-5))\] or \[y-1=\frac{3}{2}(x+5)\] yes, you got it!

OpenStudy (anonymous):

not too bad was it?

OpenStudy (anonymous):

nope! thank you so much!!!!!